the length of a rectangle is 6 ft longer than twice the width. if the perimeter is 108 ft, find the length and width of the rectangle.
Answer by Cromlix(4381)

Width = x
Length = 6 + 2x
Perimeter = 2*width + 2*length
108 = 2*x + 2* (6 + 2x)
108 = 2x + 12 + 4x
108 - 12 = 6x
96 = 6x
x = 16ft
Width = 16ft
Length = 38ft
Hope this helps.
:-)